OpenStudy (amd011802):

Graph the absolute value function f(x)=|x-2|+3

OpenStudy (mathmate):

Hint: When a function g(x) is transformed from f(x) using scaling and translation, the genera equation is: \(g(x)=a f(bx-h)+k\) where a=vertical scale factor (a>1 means stretch) 1/b=horizontal scale factor (1/b>1 means stretch) h=translate horizontally by h (h>0 means to the right) k=translate vertically by k (k>0 means up) In your given case, the parent function is |x|, h=2 k=3 Use the previous paragraph to interpret the graph of the transformed function. 

OpenStudy (amd011802):

someone please help

OpenStudy (amd011802):

ive been trying for an hour to get someone to explain this to me and help me find the answer ive posted this on so many websites.

jimthompson5910 (jim_thompson5910):

Why not use a tool like desmos? https://www.desmos.com/calculator type in the expression you want to graph into the box on the left side

jimthompson5910 (jim_thompson5910):

Here's a similar example https://www.desmos.com/calculator/ihlhmkxzl8